# The multivalued inverse of a quasiregular map is a quasiregular curve

### v1

@@ 3.1 Basic properties for branched covers
and since the sets $\overline{U}_{f}(x,2^{-j})$ are compact and nested for large enough $j$, it follows that the diameters must tend to zero. (See also [ric93, Chapter I, Lemma 4.9].)
@@
The next lemma relates the singular set of $\mathcal{A}_{d}(\Omega)$ and the branch set of $f$. Here, the singular set $\operatorname{sing}\mathcal{A}_{d}(X)$ is the union of the sets
@@
A point $y\in\Omega^{\prime}$ is in $f(B_{f})$ if and only if $f^{-1}(y)\cap B_{f}\neq\varnothing$. On the other hand, $F(y)\in\operatorname{sing}(\mathcal{A}_{d}(\Omega))$ if and only if $\iota(f,x)>1$ for some $x\in f^{-1}(y)$. These two conditions are equivalent. ∎
The following proposition will be useful later on. It is based on a theorem of Floyd [flo50], see also [lui17, Proposition 2.6], which states that compact paths admit lifts by a proper branched cover.
@@
By arguing as in [ric93, Chapter II, Corollary 3.4] using [flo50] (cf. [lui17, Proposition 2.6]), we see that $\gamma$ has $d$ total $f$-lifts $\tilde{\gamma}_{1},\ldots,\tilde{\gamma}_{d}$ which satisfy $\{j:\ \tilde{\gamma}_{j}(t)=x\}=\iota(f,x)$ for all $x\in f^{-1}(\gamma_{t})$. Relabeling these as $\gamma_{x}^{j}$ for $j=1,\ldots,\iota(f,x)$ and each $x\in f^{-1}(\gamma_{0})$, this implies $\accentset{\leftarrow}{f}(\gamma_{t})=\sum_{x\in f^{-1}(y)}\sum_{j}^{\iota(f,x)}\llbracket\gamma_{x}^{j}(t)\rrbracket$ for all $t\in[0,1]$, proving (3.1). ∎
@@
We record here the fact that the push-forward preserves continuity and measurability. This should be compared to [HKM06, Lemma 14.31] and [teripekka, Lemma 4.1] where it is shown that it preserves Sobolev regularity (resp. Lipschitz regularity when $f$ is BLD). Later on the push-forward will prove useful in connection with area formulae, see Lemma 3.6 and Proposition 3.9.
@@ 3.2 Sobolev estimates
If $\Gamma_{not}$ is the family of paths in $\Omega$ which are not locally absolutely continuous, then $\operatorname{Mod}_{n}\Gamma_{not}=0$. By Poletski’s inequality [ric93, Chapter II, Theorem 8.1] we have that $\operatorname{Mod}_{n}f(\Gamma_{not})\leq K_{I}\operatorname{Mod}_{n}\Gamma_{not}=0$. Thus, for $\operatorname{Mod}_{n}$-a.e. $\gamma$ in $\Omega^{\prime}$, every lift $\tilde{\gamma}$ of $\gamma$ is locally absolutely continuous. By Proposition 3.3 we have $\accentset{\leftarrow}{f}\circ\gamma=\llbracket\tilde{\gamma}_{1},\ldots,\tilde{\gamma}_{d}\rrbracket$, where $\tilde{\gamma}_{j}$ are the lifts of $\gamma$ by $f$ (counted with multiplicity). Thus
@@
which implies that $\accentset{\leftarrow}{f}\circ\gamma$ is absolutely continuous whenever each lift $\tilde{\gamma}_{l}$ is absolutely continuous. Thus $\accentset{\leftarrow}{f}\circ\gamma$ is absolutely continuous for $\operatorname{Mod}_{n}$-a.e. $\gamma$.
We now prove (3.3). Let $\Gamma_{0}$ be a path family in $\Omega$ with $\operatorname{Mod}_{n}\Gamma_{0}=0=\operatorname{Mod}_{n}f(\Gamma_{0})$ so that $\|Df\|$ is an upper gradient of $f$ along every $\gamma\notin\Gamma_{0}$. Observe that $f(B_{f})$ is closed since $f$ has finite degree and that $|f(B_{f})|=0$ by [ric93, Chapter II, Theorem 7.4]. If $B=B(y,r)\subset\Omega^{\prime}\setminus f(B_{f})$ is a ball such that $f^{-1}B=U_{1}\cup\cdots\cup U_{d}$, $U_{l}$ and $f:U_{l}\to B$ is quasiconformal, denote by $g_{l}:B\to U_{l}$ the inverse of $f|_{U_{l}}$, for $l=1,\ldots,d$. Then
@@
a.e. $y\in B$, see [vai71]. By Vitali’s covering theorem there exist countably many disjoint balls $B_{j}=B(y_{j},r_{j})$ as above, with inverses $g^{j}_{l}:B_{j}\to U_{l}^{j}$ satisfying (3.5) outside a null set $N_{j}\subset B_{j}$, such that $\big|[\Omega^{\prime}\setminus f(B_{f})]\setminus\bigcup_{j}B_{j}\big|=0$. Denote $E=f(B_{f})\cup\Omega^{\prime}\setminus\bigcup_{j}B_{j}$, so that $|E|=0$. For each $j$, let $\Gamma^{j}$ be a path-family in $B_{j}$ with $\operatorname{Mod}_{n}\Gamma^{j}=0$ such that $\|Dg_{l}^{j}\|$ is an upper gradient along $\gamma$, and $|\gamma^{-1}(N_{j})|=0$, for every $\gamma\notin\Gamma^{j}$. Let $\Gamma_{1}$ by the family of paths in $\Omega^{\prime}$ which have a subcurve in some $\Gamma^{j}$. Then $\operatorname{Mod}_{n}\Gamma_{1}=0$. Finally, define $\Gamma_{2}=\Gamma_{1}\cup f(\Gamma_{0})\cup\Gamma_{E}^{+}$ and note that $\operatorname{Mod}_{n}\Gamma_{2}=0$.
@@
By Theorem 3.7 (and using its notation), $(K_{I}K_{O})^{1/n}H$ is a weak upper gradient of $F$. Thus $|D\accentset{\leftarrow}{f}|\leq(K_{I}K_{O})^{1/n}H$. Using e.g. [teri-syl24, Theorem 1.1] we obtain $H\leq|D\accentset{\leftarrow}{f}|$ from the left-hand inequality in (3.3). For any compact $U\subset f(\Omega)$ we have, by Lemma 3.6, that
@@ 3.3 Area formula
Moreover, $\accentset{\leftarrow}{f}$ is metrically differentiable almost everywhere on $f(\Omega)$ and satisfies the area formula
@@
Property $(N)$ and the area formula follow from the pseudomonotonicity of $\accentset{\leftarrow}{f}$ (see Lemma 3.10 below) by results in [hei00, kar07], while the inverse Lusin property $(N^{-1})$ is a consequence of Proposition 3.4. Recall that a map $F:X\to Y$ is pseudomonotone if there exists a constants $K,r_{0}>0$ such that
@@
See [ric93, Chapter I, Lemma 4.7] for the fact that $U_{f}(x,r)$ is a normal domain of $f$ and thus $f(\partial U_{f}(x,r))=\partial B(y,r)$, implying
@@
Since $\accentset{\leftarrow}{f}\in W^{1,n}_{loc}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ is continuous and pseudomonotone, it has Lusin’s property $(N)$ by [hei00, Theorem 7.2]. Moreover, by Corollary 3.4 in [kar07] $\accentset{\leftarrow}{f}$ is metrically differentiable almost everywhere, while by Corollary 3.9(1) in [kar07], $\accentset{\leftarrow}{f}$ satisfies the area formula (3.6).
@@
4 Multivalued inverse of QR map as a QR curve
4.1 Differential forms on Almgren space
@@
4.2 Pull-back by multi-valued maps
@@
The approximate differentials $Dh_{1},\ldots,Dh_{d}$ are measurable, cf. ***. For all $x\in U$ where $h$ is approximately differentiable, we have
@@
Lemma 4.13.
The barycenter map (1.5) is $1/\sqrt{d}$-Lipschitz.
Proof.
Let $\bar{x}=\llbracket x_{1},\ldots,x_{d}\rrbracket,\bar{y}=\llbracket y_{1},\ldots,y_{d}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n})$. By relabeling we may assume that
$d_{\mathcal{A}}^{2}(\bar{x},\bar{y})=\sum_{j}^{d}|x_{j}-y_{j}|^{2}.$
Denote $z=(x_{1}-y_{1},\ldots,x_{d}-y_{d})\in(\mathbb{R}^{n})^{d}$ and recall that the average satisfies
$0\leq\sum_{j}^{d}\|z_{j}-b(z)\|^{2}=\sum_{j}^{d}\|z_{j}\|^{2}-d\|b(z)\|^{2}.$
This yields
$\|b(\bar{x})-b(\bar{y})\|^{2}=\|b(z)\|^{2}\leq\frac{1}{d}\sum_{j}^{d}\|z_{j}\|^{2}=\frac{1}{d}d_{\mathcal{A}}(\bar{x},\bar{y})^{2}.$
@@
By [del11, Proposition 1.6] there exists a ball $B_{x}=B(x,r_{x})\subset U$ and Lipschitz functions $f_{\varepsilon,x}^{i}:B_{x}\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$ ($i=0,1$) with $d_{i}<d$, $d_{0}+d_{1}=d$, $\operatorname{LIP}(f_{x}^{i})\leq CL$ and $f_{\varepsilon}=\llbracket f_{\varepsilon,x}^{0},f_{\varepsilon,x}^{1}\rrbracket$ on $B_{x}$.
@@
Note that under the assumptions in the claim $f^{*}\omega\in L^{p/k}_{loc}(U,\bigwedge^{k}\mathbb{R}^{m})$ and $f^{*}(\mathrm{d}\omega)\in L^{\frac{p}{k+1}}_{loc}(U,\bigwedge^{k+1}\mathbb{R}^{m})$ by Lemma 4.8. By considering an exhaustion of $U$ by relatively compact domains, we may assume that $U$ is bounded and that $f\in W^{1,p}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$. In particular $f\in W^{1,k}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$.
@@
4.3 Proof of Theorem 1.2
@@
$\|Dg_{j}\|^{n}\leq K_{I}\mathbf{J}g_{j}=K_{I}\star g_{j}^{*}(\mathrm{d}x).$
Observe that $\accentset{\leftarrow}{f}^{*}\omega=\sum_{j}^{d}g_{j}^{*}(\mathrm{d}x)$ and $|D\accentset{\leftarrow}{f}|^{2}=\sum_{j}^{d}\|Dg_{j}\|^{2}$ a.e. on $B(y,r)$. Consequently $\star\accentset{\leftarrow}{f}^{*}\omega=\sum_{j}^{d}\mathbf{J}g_{j}$, and thus
$|D\accentset{\leftarrow}{f}|^{n}$ $=\Big(\sum_{j}^{d}\|Dg_{j}\|^{2}\Big)^{n/2}\leq d^{n/2-1}\sum_{j}^{d}\|Dg_{j}\|^{n}$
$\leq K_{I}d^{n/2-1}\sum_{j}^{d}\mathbf{J}g_{j}=d^{n/2-1}K_{I}\star\accentset{\leftarrow}{f}^{*}\omega$
a.e. in $B(y,r)$. Since $|f(B_{f})|=0$ by [ric93, Chapter II, Proposition 4.14], the inequality holds a.e. on $f(\Omega)$ as claimed. ∎
We finish this section by providing a proof of Corollary 1.3. In fact, Corollary 1.3 follows immediately from Lemma 4.13 and the following higher integrability result for $\accentset{\leftarrow}{f}$. Indeed, the generalized inverse $g$ of $f$ is simply $g=b\circ\accentset{\leftarrow}{f}$.

### v2

@@ 3.1. Basic properties for branched covers
and since the sets $\overline{U}_{f}(x,2^{-j})$ are compact and nested for large enough $j$, it follows that the diameters must tend to zero. (See also [26, Chapter I, Lemma 4.9].)
@@
The next lemma relates the singular set of $\mathcal{A}_{d}(X)$ and the branch set of $f$. Here, the singular set $\operatorname{sing}\mathcal{A}_{d}(X)$ is the union of the sets
@@
A point $y\in f(X)$ is in $f(B_{f})$ if and only if $f^{-1}(y)\cap B_{f}\neq\varnothing$. On the other hand, $F(y)\in\operatorname{sing}(\mathcal{A}_{d}(X))$ if and only if $\iota(f,x)>1$ for some $x\in f^{-1}(y)$. These two conditions are equivalent. ∎
The following proposition will be useful later on. It is based on a theorem of Floyd [4], see also [19, Proposition 2.6], which states that compact paths admit lifts by a proper branched cover.
@@
By arguing as in [26, Chapter II, Corollary 3.4] using [4] (cf. [19, Proposition 2.6]), we see that $\gamma$ has $d$ total $f$-lifts $\tilde{\gamma}_{1},\ldots,\tilde{\gamma}_{d}$ which satisfy $\{j:\ \tilde{\gamma}_{j}(t)=x\}=\iota(f,x)$ for all $x\in f^{-1}(\gamma_{t})$. Relabeling these as $\gamma_{x}^{j}$ for $j=1,\ldots,\iota(f,x)$ and each $x\in f^{-1}(\gamma_{0})$, this implies $\accentset{\leftarrow}{f}(\gamma_{t})=\sum_{x\in f^{-1}(y)}\sum_{j}^{\iota(f,x)}\llbracket\gamma_{x}^{j}(t)\rrbracket$ for all $t\in[0,1]$, proving (3.1). ∎
@@
We record here the fact that the push-forward preserves continuity and measurability. This should be compared to [9, Lemma 14.31] and [23, Lemma 4.1] where it is shown that it preserves Sobolev regularity (resp. Lipschitz regularity when $f$ is BLD). Later on the push-forward will prove useful in connection with area formulae, see Lemma 3.6 and Proposition 3.9.
@@ 3.2. Sobolev estimates
If $\Gamma_{not}$ is the family of paths in $\Omega$ which are not locally absolutely continuous, then $\operatorname{Mod}_{n}\Gamma_{not}=0$. By Poletski’s inequality [26, Chapter II, Theorem 8.1] we have that $\operatorname{Mod}_{n}f(\Gamma_{not})\leq K_{I}\operatorname{Mod}_{n}\Gamma_{not}=0$. Thus, for $\operatorname{Mod}_{n}$-a.e. $\gamma$ in $\Omega^{\prime}$, every lift $\tilde{\gamma}$ of $\gamma$ is locally absolutely continuous. By Proposition 3.3 we have $\accentset{\leftarrow}{f}\circ\gamma=\llbracket\tilde{\gamma}_{1},\ldots,\tilde{\gamma}_{d}\rrbracket$, where $\tilde{\gamma}_{j}$ are the lifts of $\gamma$ by $f$ (counted with multiplicity). It follows that
@@
which implies that $\accentset{\leftarrow}{f}\circ\gamma$ is absolutely continuous whenever each lift $\tilde{\gamma}_{l}$ is absolutely continuous. In particular, $\accentset{\leftarrow}{f}\circ\gamma$ is absolutely continuous for $\operatorname{Mod}_{n}$-a.e. $\gamma$.
We now prove (3.3). Let $\Gamma_{0}$ be a path family in $\Omega$ with $\operatorname{Mod}_{n}\Gamma_{0}=0=\operatorname{Mod}_{n}f(\Gamma_{0})$ so that $\|Df\|$ is an upper gradient of $f$ along every $\gamma\notin\Gamma_{0}$. Observe that $f(B_{f})$ is closed since $f$ has finite degree and that $|f(B_{f})|=0$ by [26, Chapter II, Theorem 7.4]. If $B=B(y,r)\subset\Omega^{\prime}\setminus f(B_{f})$ is a ball such that $f^{-1}B=U_{1}\cup\cdots\cup U_{d}$, $U_{l}$ and $f:U_{l}\to B$ is quasiconformal, denote by $g_{l}:B\to U_{l}$ the inverse of $f|_{U_{l}}$, for $l=1,\ldots,d$. Then
@@
a.e. $y\in B$, see [29]. By Vitali’s covering theorem there exist countably many disjoint balls $B_{j}=B(y_{j},r_{j})$ as above, with inverses $g^{j}_{l}:B_{j}\to U_{l}^{j}$ satisfying (3.5) outside a null set $N_{j}\subset B_{j}$, such that $\big|[\Omega^{\prime}\setminus f(B_{f})]\setminus\bigcup_{j}B_{j}\big|=0$. Denote $E=f(B_{f})\cup\Omega^{\prime}\setminus\bigcup_{j}B_{j}$, so that $|E|=0$. For each $j$, let $\Gamma^{j}$ be a path-family in $B_{j}$ with $\operatorname{Mod}_{n}\Gamma^{j}=0$ such that $\|Dg_{l}^{j}\|$ is an upper gradient along $\gamma$, and $|\gamma^{-1}(N_{j})|=0$, for every $\gamma\notin\Gamma^{j}$. Let $\Gamma_{1}$ by the family of paths in $\Omega^{\prime}$ which have a subcurve in some $\Gamma^{j}$. Then $\operatorname{Mod}_{n}\Gamma_{1}=0$. Finally, define $\Gamma_{2}=\Gamma_{1}\cup f(\Gamma_{0})\cup\Gamma_{E}^{+}$ and note that $\operatorname{Mod}_{n}\Gamma_{2}=0$.
@@
By Theorem 3.7 (and using its notation), $(K_{I}K_{O})^{1/n}H$ is a weak upper gradient of $F$. Thus $|D\accentset{\leftarrow}{f}|\leq(K_{I}K_{O})^{1/n}H$. Using e.g. [3, Theorem 1.1] we obtain $H\leq|D\accentset{\leftarrow}{f}|$ from the left-hand inequality in (3.3). For any compact $U\subset f(\Omega)$ we have, by Lemma 3.6, that
@@ 3.3. Area formula
Moreover, $\accentset{\leftarrow}{f}$ is differentiable almost everywhere on $f(\Omega)$ and satisfies the area formula
@@
Property $(N)$ and the area formula follow from the pseudomonotonicity of $\accentset{\leftarrow}{f}$ (see Lemma 3.10 below) by results in [10, 17], while the inverse Lusin property $(N^{-1})$ is a consequence of Proposition 3.4. Recall that a map $F:X\to Y$ is pseudomonotone if there exists a constants $K,r_{0}>0$ such that
@@
See [26, Chapter I, Lemma 4.7] for the fact that $U_{f}(x,r)$ is a normal domain of $f$ and thus $f(\partial U_{f}(x,r))=\partial B(y,r)$, implying
@@
Since $\accentset{\leftarrow}{f}\in W^{1,n}_{loc}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ is continuous and pseudomonotone, it has Lusin’s property $(N)$ by [10, Theorem 7.2]. Moreover, by Corollary 3.4 in [17] $\accentset{\leftarrow}{f}$ is metrically differentiable almost everywhere, while by Corollary 3.9(1) in [17], $\accentset{\leftarrow}{f}$ satisfies the area formula (3.6).
The a.e. metric differentiability of $f$ implies that $\operatorname{Lip}f(x)<\infty$ a.e. $x\in f(\Omega)$, thus by the Stepanov theorem for multivalued maps (see [20, Corollary 2.8]) it follows that $f$ is differentiable almost everywhere.
@@
4. Pull-back of differential forms by multi-valued Sobolev maps
4.1. Differential forms on Almgren space
@@
4.2. Pull-back by multi-valued maps
@@
The approximate differentials $Dh_{1},\ldots,Dh_{d}$ are measurable. For all $x\in U$ where $h$ is approximately differentiable, we have
@@
By [2, Proposition 1.6] there exists a ball $B_{x}=B(x,r_{x})\subset U$ and Lipschitz functions $f_{\varepsilon,x}^{i}:B_{x}\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$ ($i=0,1$) with $d_{i}<d$, $d_{0}+d_{1}=d$, $\operatorname{LIP}(f_{x}^{i})\leq CL$ and $f_{\varepsilon}=\llbracket f_{\varepsilon,x}^{0},f_{\varepsilon,x}^{1}\rrbracket$ on $B_{x}$.
@@
Note that under the assumptions in the claim, we have $f^{*}\omega\in L^{p/k}_{loc}(U,\bigwedge^{k}\mathbb{R}^{m})$ and $f^{*}(\mathrm{d}\omega)\in L^{\frac{p}{k+1}}_{loc}(U,\bigwedge^{k+1}\mathbb{R}^{m})$ by Lemma 4.8. By considering an exhaustion of $U$ by relatively compact domains, we may assume that $U$ is bounded and that $f\in W^{1,p}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$. In particular $f\in W^{1,k}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$.
@@
5. Multivalued inverse of QR map as a QR curve
Throughout this section, $\omega=\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}})$ denotes the canonical $n$-form on $\mathcal{A}_{d}(\mathbb{R}^{n})$.
Lemma 5.1.
Let $f\in W^{1,n}_{loc}(\Omega,\mathcal{A}_{d}(\mathbb{R}^{n}))$, where $\Omega$ is an open set. Then
$\star f^{*}\omega=\sum_{j}^{d}\det(Df_{j})\quad\textrm{ a.e. on }\Omega.$ (5.1)
where $Df=\llbracket Df_{1},\ldots,Df_{d}\rrbracket$ is the weak differential of $f$. In particular, we have the inequality
$|\star f^{*}\omega|\leq\mathbf{J}f\quad\textrm{a.e. on }U.$ (5.2)
Proof.
Let $Df=\llbracket Df_{1}(x),\ldots,Df_{d}(x)\rrbracket$ be the differential of $f$ at $x$. By Definitions 4.4 and 4.6 we have that
$(f^{*}\omega)_{x}(v_{1},\ldots,v_{n})$ $=\omega_{f(x)}(Df(x)v_{1},\ldots,Df(x)v_{n})$
$=\sum_{j}^{d}Df_{j}(x)v_{1}\wedge\cdots\wedge Df_{j}(x)v_{n}$
$=\sum_{j}^{d}\det(Df_{j}(x))v_{1}\wedge\cdots\wedge v_{n},\quad v_{1},\ldots,v_{n}\in T_{x}U,$
which implies (5.1). The estimate (5.2) follows from (5.1) together with Lemma 2.4 and Minkowski’s determinant inequality:
$|\star f^{*}\omega|$ $\leq\sum_{j}^{d}|\det(Df_{j})|=\sum_{j}^{d}[\det(Df_{j}^{T}Df_{j})]^{1/2}$
$\leq\Big(\sum_{j}^{d}[\det(Df_{j}^{T}Df_{j})]^{1/n}\Big)^{n/2}\leq\det\Big(\sum_{j}^{d}Df_{j}^{T}Df_{j}\Big)^{1/n\ \cdot\ n/2}=\mathbf{J}f.$
@@ 5. Multivalued inverse of QR map as a QR curve
$\|Dg_{j}\|^{n}\leq K_{I}J_{g_{j}}=K_{I}\star g_{j}^{*}(\mathrm{d}x).$
Observe that $\star\accentset{\leftarrow}{f}^{*}\omega=\sum_{j}^{d}J_{g_{j}}$ and $|D\accentset{\leftarrow}{f}|^{2}\leq\sum_{j}^{d}\|Dg_{j}\|^{2}$ a.e. on $B(y,r)$. Thus
$|D\accentset{\leftarrow}{f}|^{n}$ $\leq\Big(\sum_{j}^{d}\|Dg_{j}\|^{2}\Big)^{n/2}\leq d^{n/2-1}\sum_{j}^{d}\|Dg_{j}\|^{n}$
$\leq K_{I}d^{n/2-1}\sum_{j}^{d}J_{g_{j}}=d^{n/2-1}K_{I}\star\accentset{\leftarrow}{f}^{*}\omega$
a.e. in $B(y,r)$. Since $|f(B_{f})|=0$ by [26, Chapter II, Proposition 4.14], the inequality holds a.e. on $f(\Omega)$ as claimed. ∎
Remark 5.2.
An inspection of the proof yields the following inequality:
$\mathbf{J}\accentset{\leftarrow}{f}\leq K_{I}\star\accentset{\leftarrow}{f}^{*}\omega\quad\textrm{ a.e. on }f(\Omega).$ (5.3)
This a reverse version (up to the multiplicative constant $K_{I}$) of the inequality in Lemma 5.1. Unlike Lemma 5.1, (5.3) does not hold for general Sobolev maps.
The following higher integrability result for $\accentset{\leftarrow}{f}$ implies part of Corollary 1.3.